OpenStudy (anonymous):

Alan saved $50,000 for 10 tens years. He earned $25,000 interest from his savings account. what was his bank interest rate: A. 20% B. 200% C. 5% D. 0.05%

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

is this simple or compound interest?

OpenStudy (anonymous):

I'm not sure it doesn't say....

OpenStudy (anonymous):

I'll assume it's compound. :) When the interest is compounded once a year: A = P(1 + r)n Where: P is the principal (the initial amount you borrow or deposit) r is the annual rate of interest (percentage) n is the number of years the amount is deposited or borrowed for. A is the amount of money accumulated after n years, including interest. So... (50000+25000)=50000(1+r)^10 75000=50000(1+r)^10 1.5 = (1+r)^10 (1.5)^(1/10)=1+r Solving for r, we get r=0.041379... so the rate is roughly 4.14% :)

OpenStudy (anonymous):

Ok thanks